OpenStudy (anonymous):

(3times square root of 5 + square root of 3) (square root of 5 -7 square root of 3)

OpenStudy (anonymous):

\[(3\sqrt{5}+\sqrt{3}) (\sqrt{5}-7\sqrt{3})\]